1. Solve by making a substitution to reduce the given second order De to a first order DE. 1. x? y + 2xy - 1 = 0, x>0 (ans.
Answer #1

bolly! + 2xy-1=0, x70 y=v ixy=v det, x -Utsul 12 - y = 0 16)+ 4.0 . Depy = -4+x60 putting values from & & 6 in givens egrsay = (t dx- & Site de - y loga ( + ((s: constant) => y = logx + 43+ y= 4x + co + lux Q y+ y (y)² = 0 let y=4 ulou y- du312 g Sy-26) dy = $20x y3 - 244 + 200 3 3 yy=2 y(0)=1 Y(0) = 2 det y=2 - y=2 - 22=2 2d2= 2dx untegrating both side, -> =